# Generative Motion Stylization within Canonical Motion Space

### v1

@@ 4.2. Quantitative Results
Content preservation. As shown in Table 1, Motion effectively preserves the source motion content compared to the baselines. Despite SinMDM achieving the best FMD, its generated results do not express the motion style, as illustrated in Figure 5. The results of DME and Motion Puzzle reveal lower generalization ability than Motion, as their performance on Xia source motion is significantly better than on SMPL source motion. However, on both datasets, their performance is worse than Motion. Additionally, Motion (CLIP) and (w/o TET) exhibit worse performance on this metric than ours, further validating the effectiveness of our key designs.
@@
Stylized motion diversity. Thanks to the diffusion model we employed, Motion is capable of generating diverse motion from a single source sequence. The Global Diversity and Local Diversity are calculated based on the source motion. SinMDM obtains the highest values since its generated result is similar to the source motion while lacking distinct style properties. In contrast, the results of SinMDM are primarily based on the style motion, leading to a deficiency in motion content. On the other hand, the results of Motion preserve the source motion content, express style motion features, and exhibit rich diversity.
@@ 4.3. Additional Evaluation
We evaluated Motion on the following two applications:
Manual adjustment. As discussed in Section 3.3, the network flow of TSD in our Motion is bifurcated into two branches to ensure stable training. During inference, only the stylization branch is utilized for generating the stylized motion. To enhance the flexibility of the pipeline, we introduce a weight parameter , to blend the two branches and perform feature linear interpolation in the canonical motion space. This allows for manual adjustment, providing control over the degree of motion stylization. Figure 10 illustrates the cases of manually adjusting . By gradually scaling , we can obtain results that transition smoothly from the source motion to the stylized motion, enabling interactive selection of visually optimal results.
@@ 5. Conclusions
In this work, we propose a novel generative motion stylization pipeline, Motion, capable of synthesizing diverse and stylized motion from a single source sequence using multi-modality style descriptions. In Motion, two key strategies are exploited to embed the cross-modality style prompts and the cross-structure skeleton motion into a canonical motion space. The first strategy is cross-modality style embedding, aligning style motion with the CLIP space and extracting style features from motion, text, or image prompts. The second strategy is cross-structure topology shifting, aligning arbitrary skeleton structures with the SMPL skeleton in latent space, enabling general motion stylization for various characters. In implementation, we construct Motion based on the topology-shifted stylization diffusion model and pre-train the style embedding auto-encoder on a large-scale motion, text, and image triplet dataset. Subsequently, we transfer the pre-trained weights of the style decoder into the diffusion model to ensure the reconstruction of the dynamic motion style. We conduct extensive experiments to validate the effectiveness of our method, demonstrating that it achieves state-of-the-art performance compared to baseline methods.
Limitations. One potential drawback lies in the limitation of style descriptions that Motion trained on, possibly leading to the generated results that may not accurately express the motion style described in arbitrary custom prompts. Furthermore, foot contact is not our primary focus, it can be addressed using the method proposed in (Aberman et al., 2020a).

### v2

@@ 4.2. Quantitative Results
Content preservation. As shown in Table 1, Motion effectively preserves the source motion content compared to the baselines. Despite SinMDM∗ achieving the best FMD, its generated results do not express the motion style, as illustrated in Figure 5. The results of DME and Motion Puzzle reveal lower generalization ability than Motion, as their performance on Xia source motion is significantly better than on SMPL source motion. However, on both datasets, their performance is worse than Motion. Additionally, Motion (CLIP) and (w/o TET) exhibit worse performance on this metric than ours, further validating the effectiveness of our key designs.
@@
Stylized motion diversity. Thanks to the diffusion model we employed, Motion is capable of generating diverse motion from a single source sequence. The Global Diversity and Local Diversity are calculated based on the source motion. SinMDM obtains the highest values since its generated result is similar to the source motion while lacking distinct style properties. In contrast, the results of SinMDM∗ are primarily based on the style motion, leading to a deficiency in motion content. On the other hand, the results of Motion preserve the source motion content, express style motion features, and exhibit rich diversity.
@@ 4.3. Additional Evaluation
Manual adjustment. As discussed in Section 3.3, the network flow of TSD in our Motion is bifurcated into two branches to ensure stable training. Only the stylization branch is utilized during inference to generate the stylized motion. To enhance the flexibility of the pipeline, we introduce a weight parameter , to blend the two branches and perform feature linear interpolation in the canonical motion space. This allows for manual adjustment, providing smooth control over the degree of motion stylization. By gradually scaling , as Figure 10 illustrates, we can obtain results that transition smoothly from the source motion to the stylized motion, enabling interactive selection of visually optimal results.
@@ 5. Conclusions
In this work, we propose a novel generative motion stylization pipeline, Motion, capable of synthesizing diverse and stylized motion from a single source sequence using multi-modality style descriptions. In Motion, two key strategies are exploited to embed the cross-modality style prompts and the cross-structure skeleton motion into a canonical motion space. The first strategy is cross-modality style embedding, aligning style motion with the CLIP space and extracting style features from motion, text, or image prompts. The second strategy is cross-structure topology shifting, aligning arbitrary skeleton structures with the SMPL skeleton in latent space, enabling general motion stylization for various characters. In implementation, we construct Motion based on the topology-shifted stylization diffusion model and pre-train the style embedding auto-encoder on a large-scale motion, text, and image triplet dataset. Subsequently, we transfer the pre-trained weights of the canonical decoder into the diffusion model to ensure the reconstruction of the dynamic motion style. We conduct extensive experiments to validate the effectiveness of our method, demonstrating that it achieves the state-of-the-art performance compared to the baseline methods.
Limitations. One potential drawback lies in the limitation of style descriptions that Motion trained on, possibly leading to the generated results that may not accurately express the motion style described in arbitrary custom prompts. Furthermore, foot contact is not our primary focus, it can be addressed using the method proposed in (Aberman et al., 2020a). We acknowledge that some of the generated motion of our Motion exhibits motion distortions and artifacts. However, it should be noted that stylizing motion on cross-structure characters using cross-modality sources remains a challenging task that has not been completely solved. We are committed to continuing our efforts to improve the geometric quality of generated motion.
Acknowledgements. This work was supported by the Natural Science Fund for Distinguished Young Scholars of Hubei Province No.2022CFA075, the National Natural Science Foundation of China (NSFC) No.62106177, the Fundamental Research Funds for Central Universities No.2042023KF0180.